Tigist wishes to mix two types of food C and D in such a way that the vitamin contents of the mixture contain at least 8 units of vitamin C and 11 units of vitamin D. Food C costs $ 60/kg and Food D costs $80/kg. Food C contains 3 units/kg of Vitamin A and 5 units / kg of Vitamin B while food D contains 4 units/kg of Vitamin A and 2 units/kg of vitamin B. Determine the minimum cost of the mixture by graphic model.



1
Expert's answer
2022-04-04T16:04:25-0400

"x_1-mass\\,\\,of\\,\\,food\\,\\,C\\\\x_2-mass\\,\\,of\\,\\,food\\,\\,D\\\\\\left\\{ \\begin{array}{c}\t3x_1+4x_2\\geqslant 8\\\\\t5x_1+2x_2\\geqslant 11\\\\\\end{array} \\right. \\\\Z=60x_1+80x_2\\rightarrow \\min \\\\We\\,\\,see\\,\\,from\\,\\,the\\,\\,graph\\,\\,that\\,\\,\\min imum\\,\\,is\\,\\,for\\\\\\left\\{ \\begin{array}{c}\t3x_1+4x_2=8\\\\\t5x_1+2x_2=11\\\\\\end{array} \\right. \\Rightarrow x_1=2,x_2=0.5\\\\Z_{\\min}=60\\cdot 2+80\\cdot 0.5=160"